OpenStudy (anonymous):

2x^2+3x=11 using quadratic formula

OpenStudy (imstuck):

First things first...do you know the quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (imstuck):

Ok, then secondly, you need that polynomial in standard form. That means that you need to move the 11 over to be with thee rest of the equation.

OpenStudy (imstuck):

\[2x ^{2}+3x-11=0\]

OpenStudy (anonymous):

okay i have that written down. what is next ?

OpenStudy (imstuck):

Here is the formula for a = 2, b = 3, c = -11:

OpenStudy (imstuck):

\[\frac{ -3\pm \sqrt{(3)^{2}-4(2)(-11)} }{ 2(2) }\]can you do that math?

OpenStudy (imstuck):

do inside the radical sign first...

OpenStudy (anonymous):

okay i did and i got 88+9

OpenStudy (imstuck):

which is what? what does that equal?

OpenStudy (anonymous):

equals 97

OpenStudy (anonymous):

my answer choices are 0 1 and 17

OpenStudy (imstuck):

That's right and is 97 have a perfect square?

OpenStudy (imstuck):

we are not to the answer stage just yet...

OpenStudy (anonymous):

no it doesnt

OpenStudy (imstuck):

so we have this then:\[\frac{ -3\pm \sqrt{97} }{ 4 }\]and if your problem was typed here correctly, that is the answer. It is not 0, 1 or 17.

OpenStudy (imstuck):

let me check myself right quick...

OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? 2x2 + 3x = -1

OpenStudy (anonymous):

this is the problem

OpenStudy (imstuck):

you are looking for the discriminant of 2x^2, + 3x = -1?

OpenStudy (anonymous):

if that ios the value of b2-4ac than yes

OpenStudy (imstuck):

I thought you were to factor the problem with the -11 in it? which is it? finding b^2-4ac is the discriminant.

OpenStudy (imstuck):

Ok, redo.

OpenStudy (imstuck):

This is the problem...yes?\[2x ^{2}+3x=-1\]

OpenStudy (imstuck):

a = 2, b = 3, c = +1 (when you move it over to put the equation into standard form)

OpenStudy (imstuck):

The discriminant is \[b ^{2}-4ac\]which is, using your values,\[(3)^{2}-4(2)(1)\]That equals 1.

OpenStudy (imstuck):

your answer is 1. It is there in your choices.